An adjustable monochromatic laser beam is directed at two slits spaced 0.400 mm apart. A screen 3 m away displays the interference pattern due to the diffraction of the beam by the two slits. If the laser wavelength is changed uniformly from 500 nm to 630 nm over 5 s, what is the speed of the first maximum intensity line from the center? Is it moving toward the center or away from the center?
1
Expert's answer
2020-11-16T07:49:40-0500

As per the given question,

Distance between the two slits "(d)=0.4mm = 4\times 10^{-4}m"

Distance between the slits and the screen "(D)=3m"

Initial wavelength "(\lambda_i)=500nm = 5\times 10^{-7}m"

Final wavelength "(\lambda_f) =630nm = 6.3\times 10^{-7}m"

Time (t)=3 sec

distance of the first maximum central intensity "(y_i) = \frac{n \lambda_i D}{d}"

Now, substituting the values, "(y_i)=\frac{ 5\times 10^{-7}m\times 3}{4\times 10^{-4}}m"

"\Rightarrow y_i=3.75\times 10^{-3}m"

"\Rightarrow y_i =3.65mm"

Similarly,"y_f=\frac{6.3\times 10^{-7}\times3}{4\times 10^{-4}}m"

"y_f=4.724mm"

Hence, the speed of the first maximum intensity "=\frac{4.724-3.65}{3}m/s"

"=0.358m\/s"

It is moving away from the center, because the wavelength is increasing.